Suppose the demand for a product is given by Qc=200 - 2Pt+Pc where Qc is the quantity demand for coffee,Pc is the price of coffee,Pt is the price of the substitute good tea.The price of the substitute good is  ₹400/kg.Suppos Pc=₹620/kg what would be the demand of coffee?


1
Expert's answer
2021-02-22T14:00:01-0500

Let the demand for a product is given by


"Q_c=200-2P_t+P_c."

If the price of the substitute good tea is ₹400/kg and the price of coffee is ₹620/kg, the demand of coffee would be:


"Q_c=200-2\cdot400+620=20\ units."